The instructions say to solve by factoring. x²+13x=-42. There are supposed to be two answers. I know how to factor when the equation equals 0, but not when there is an answer other than 0.
Answer by ewatrrr(24785)

 
Hi,
Make IT equal zero
x^2 + 13x + 42 = 0
(x+6)(x+7) = 0
(x+6)= 0, x = -6
(x+7) = 0, x = -7